# From tug-of-war to Brownian Boost: explicit ODE solutions for player-funded stochastic-differential games

### v3

@@ 1.7.2. Incentive Inch, Outcome Mile
$\lambda_{\rm max}(0,\rho)\,=\,\sup\,\left\{\,\frac{\int_{\mathbb{R}}g_{\rho}(x,u)\,{\rm d}u}{\int_{\mathbb{R}}f_{\rho}(x,u)\,{\rm d}u}:x\in\mathbb{R}\,\right\}$ (17)
@@
As we will see in Proposition 5.4(5), the $f$- and $g$-integrals are always equal, so $\lambda_{\rm max}(0,\rho)=1$. This holds for any $\rho\in(0,\infty)$, though the interpretation via Nash equilibria is known for $\mathrm{BB}(\rho)$ only when $\rho\in(0,1]$, via $\mathrm{TLP}(\kappa,\rho)$ and Theorem 1.8.
To interpret this conclusion, we review the discouragement effect. Suppose that in a game of $\mathrm{BB}(\rho)$, $\int_{\mathbb{R}}g_{\rho}(x,u)\,{\rm d}u>\int_{\mathbb{R}}f_{\rho}(x,u)\,{\rm d}u$. Mina has a greater incentive and, we may speculate, would be prepared to out-stake Maxine by a constant factor at every instant in the game. A constant negative drift would result and Mina would win. Maxine will recognize this at the outset, become discouraged, stake nothing, and permit Mina to win at arbitrarily small running cost.
This heuristic should hardly be readily accepted, but it is coherent with $\lambda_{\rm max}(0,\rho)=1$ and the non-existence of equilibria in the imbalanced game. (That said, the argument suggests that the more incentivized player will win $\mathrm{BB}(\rho)$ at no running cost. But the absence of equilibria in the game gives neither this player nor her opponent any guidance as to how to play it.)
@@
There exist $C>0$ and $c,\kappa_{0}\in(0,1)$ such that, for $\kappa\in(0,\kappa_{0})$ and $\rho\in(0,1]$, $\big|\lambda_{\rm max}(\kappa,\rho)-1\big|\leq C\kappa^{c}$.
Remark. The result may be extended to the regime $\rho>1$ when $(\kappa,\rho)\in W$ (that is, $\kappa^{2}\rho\leq 1$), with $c=c(\rho)$ decaying to zero in the high-$\rho$ limit.
In the final Section 7, we report on $\lambda_{\rm max}:(0,1]^{2}\to[1,\infty)$ numerically, finding this function to have some remarkable features. The numerics prompt the following conjecture.
@@ 1.7.3. Structure of the article
There are six further sections. In Section 2, we introduce several basic elements that undergird our analysis of ABMN($\kappa,\rho$) elements and their game-theoretic significance including a solution of the one-step game. In Section 3, we show that $s$ acts as $s(\phi_{i})=\phi_{i+1}$ and prove the ABMN($\kappa,\rho$) explicit form Theorem 1.16. Developing $s$-orbit asymptotics, we then prove the fixed-parameter ABMN($\kappa,\rho$) Theorem 1.21. The Nash-ABMN equivalence Theorem 1.8 is proved in Section 4. We then turn to Brownian Boost, showing heuristically that its equilibria are governed by solutions of the $\mathrm{BB}(\rho)$ ODE pair, and giving an analytic study of these solutions in Section 5. In Section 6, we represent Brownian Boost as a high-noise limit of the Trail of Lost Pennies and prove the low-$\kappa$ $\lambda_{\rm max}$ Theorem 1.22 and the asymptotic stakes-and-gameplay Theorem 1.18. The final Section 7 is devoted to presenting numerical findings (including striking behaviour for the map $\lambda_{\rm max}$) and several directions for further inquiry.
@@ 1.7.4. Acknowledgements
The author warmly thanks a referee of [26] whose pivotal role has been indicated in Section 1.2. He is supported by NSF-DMS grants $2153359$ and $2450726$.
2. Some basic symmetries and tools
We introduce several basic tools for analysing elements of ABMN($\kappa,\rho$). In consecutive subsections, a role-reversal symmetry is used to invert the map $s(\phi_{0})=\phi_{1}$; the escape of forward and backward orbits of $s$ is noted; the battlefield index of an element of ABMN($\kappa,\rho$) is specified via the $s$-orbit; and a useful device for studying such elements, the Mina margin map, is defined. In a final subsection, we analyse Penny Forfeit, the one-step sub-game of $\mathrm{TLP}(\kappa,\rho)$.
2.1. Role-reversal symmetry and the inverse of $s$
The solution class ABMN($\kappa,\rho$) is invariant under $\mathbb{Z}$-shift. A further role-reversal symmetry gives a formula for the inverse of $s$.
Proposition 2.1.
The function $s:(0,\infty)\to(0,\infty)$ from Definition 1.13(1) is invertible, with s1(x)= $s^{-1}(x)\,=\,1$/s(1/x) for  $x\in(0,\infty)$.
Proof. By definition, $s$ sends $\phi_{0}$ to $\phi_{1}$. Since both maps $\phi_{i}(\beta)=\phi_{i}(\kappa,\rho,\beta)$ are bijections $(0,\infty)\to(0,\infty)$, the inverse map $s^{-1}$ sending $\phi_{1}$ to $\phi_{0}$ is well defined. The formula s1(x)= $s^{-1}(x)\,=\,1$/s(1/x) amounts to $1/\phi_{0}=s(1/\phi_{1})$. To see this, note from (10) and (11) that when $\phi_{0}=\phi_{0}(\kappa,\rho,\beta)$ and $\phi_{1}=\phi_{1}(\kappa,\rho,\beta)$, we have that $1/\phi_{0}=\phi_{1}(\kappa,\rho,1/\beta)$ and $1/\phi_{1}=\phi_{0}(\kappa,\rho,1/\beta)$. So the sought equality $1/\phi_{0}=s(1/\phi_{1})$ is then the instance of $s(\phi_{0})=\phi_{1}$ corresponding to $1/\beta$. ∎
A game-theoretic view of the symmetry underlying the preceding argument may help to elucidate its opaque algebraic satisfaction of the needed condition. For an $(a,b,m,n)\in$ ABMN($\kappa,\rho$), $\beta=\beta_{0}$ given by $n_{1,-1}/m_{-1,1}$ from Definition 3.2 parameterises $\phi_{0}=n_{0,-1}/m_{-1,0}$ and $\phi_{1}=n_{1,0}/m_{0,1}$. Reflect gameplay governed by the $(a,b)$ strategy pair through the origin. The players now stand at the wrong ends, and under the reflected gameplay each would play against her own interest. But their play makes sense if they now change ends. The new gameplay is governed by the strategy pair $\big(b(-\,\bullet),a(-\,\bullet)\big)$. This pair extends to $\big(b(-\,\bullet),a(-\,\bullet),n(-\,\bullet),m(-\,\bullet)\big)\in$ ABMN($\kappa,\rho$). The switch from old to new solution maps $\phi_{0}\mapsto 1/\phi_{1}$, $\phi_{1}\mapsto 1/\phi_{0}$, and $\beta_{0}\mapsto 1/\beta_{0}$, which explains the relation s1(x)= $s^{-1}(x)\,=\,1$/s(1/x) as well as the reciprocal $\beta$-parametrization that appears in the proof.
Corollary 2.2.
If $(a,b,m,n)$ is an element of ABMN($\kappa,\rho$), then so is
@@ 2.2. The bijections $\phi_{0}$ and $\phi_{1}$, and the orbit of $s$
The next result records some basic properties of $\phi_{0}$ and $\phi_{1}$ which have permitted the specification of the map $s:(0,\infty)\to(0,\infty)$ that sends $\phi_{0}$ to $\phi_{1}$ in Definition 1.13.
@@
Figure 2.1. The main contours on the $(\kappa,\rho)$-map. The curve $C_{1}$ is the locus of $\kappa^{2}\rho=1$, which is the upper boundary of the region $W$ in (8); $C_{2}$ is the locus of $\kappa\rho=1+\sqrt{1-\kappa^{2}}$. The curves emanate from $(1,1)$, the point indexing the game studied in [26] and the upper-right corner of the unit box in which Nash-ABMN equivalence is established by Theorem 1.8. The $\mathrm{BB}(\rho)$-line $\kappa=0$ lies below $C_{1}$, which is indicative of how the ODE-pair Theorem 1.3 is valid for all $\rho\in(0,\infty)$.
@@
Writing $q$ for the right-hand denominator in (10), $\frac{\partial\,\phi_{0}(\kappa,\rho,\beta)}{\partial\beta}\,=\,P(\beta^{\rho})q^{-2}$ where of the coefficients of the quartic $P(x)=\sum_{i=0}^{4}h_{i}x^{i}$, $h_{0}=(1+\kappa)^{2}$ and $h_{4}=(1-\kappa)^{2}$ are evidently non-negative. That $h_{2}=2\big(3-\kappa^{2}(1+2\rho^{2})\big)$ is likewise follows from the conditions, which are weaker than our hypothesis, that $|\kappa|\leq 1$ and $\rho|\kappa|\leq 1$. The coefficients $h_{1}=4(1-\kappa)(1-\kappa\rho^{2})$ and $h_{3}=4(\kappa+1)(\rho^{2}\kappa+1)$ are also non-negative: in one case trivially; in the other, as our hypothesis is tailored to show; and with the sign of $\kappa$ determining which case applies.
@@
where $g(\kappa,x)=(1-\kappa)x^{2}+2(1-\rho\kappa)x+1+\kappa$. When $\kappa\in(0,1)$ and $\rho>0$, the quadratic $g(-\kappa,\cdot):\mathbb{R}\to\mathbb{R}$ is always positive, while for such $\kappa$-values, $g(\kappa,\cdot):\mathbb{R}\to\mathbb{R}$ is as well, because the discriminant sign condition $\rho<\kappa^{-1}(1+\sqrt{1-\kappa^{2}})$ has been hypothesised. Thus, $\phi_{0}$ exceeds $\phi_{1}$ when $\kappa\in(0,1)$.
@@ 2.3. The battlefield index
Proof: (1). The function $s=s(\kappa,\rho,\cdot):(0,\infty)\to(0,\infty)$ belongs to $\rm IB$}, so it meets the decreasing map $1/x$ at exactly one $x\in(0,\infty)$.
@@ 2.4. The Mina margin map
In Definition 1.10, $\lambda_{\rm max}(\kappa,\rho)$ has been defined to be the supremum of the Mina margin $n_{\infty,-\infty}/m_{-\infty,\infty}$ over all ABMN($\kappa,\rho$) solutions. It is worth noting that the several symmetries enjoyed by ABMN($\kappa,\rho$) permit a more restricted supremum to be taken, and the Mina margin map is a useful device for making this point. Recall from Section 1.5 that for $x\in(0,\infty)$ there is a unique element of ABMN($\kappa,\rho$) with $m_{-\infty}=n_{\infty}=0$, $m_{\infty}=1$ and $\frac{n_{0,-1}}{m_{-1,0}}=x$. This is the standard solution $\big(a^{\rm st}_{i}(x),b^{\rm st}_{i}(x),m^{\rm st}_{i}(x),n^{\rm st}_{i}(x):i\in\mathbb{Z}\big)$.
@@
$\mathcal{M}_{\kappa,\rho}^{-k,k}(x)\,\longrightarrow\,\mathcal{M}_{\kappa,\rho}(x)\,$, (18)
@@
As $k\to\infty$, the left-hand side converges to $\mathcal{M}_{\kappa,\rho}\big(s(x)\big)$, by (18) with $x\to s(x)$, while the right-hand side converges to $\frac{n^{\rm st}_{\infty,-\infty}}{m^{\rm st}_{-\infty,\infty}}=\mathcal{M}_{\kappa,\rho}(x)$ by (18) and the decay of high-indexed $m$- and $n$-differences in Theorem 1.21. Hence $\mathcal{M}_{\kappa,\rho}(s(x))=\mathcal{M}_{\kappa,\rho}(x)$ for $x\in(0,\infty)$.
@@
(2). The finite-trail Mina margin map may be expressed as a ratio

### v4

@@ 1.5.3. Incentive Inch, Outcome Mile
$\lambda_{\rm max}(0,\rho)\,=\,\sup\,\left\{\,\frac{\int_{\mathbb{R}}g_{\rho}(\xi,r)\,{\rm d}r}{\int_{\mathbb{R}}f_{\rho}(\xi,r)\,{\rm d}r}:\xi\in(0,\infty)\,\right\}$ (1.18)
@@
We have already noted a result that is formally stated as Proposition 5.4(5): the $f$- and $g$-integrals are always equal, so $\lambda_{\rm max}(0,\rho)=1$. This holds for any $\rho\in(0,\infty)$, though the interpretation via Nash equilibria is known for $\mathrm{BB}(\rho)$ only when $\rho\in(0,1]$, via $\mathrm{TLP}(\kappa,\rho)$ and Theorem 1.3. This is the conclusion of the second preface premise, validated: the non-existence of equilibria in the imbalanced $\lambda\neq 1$ game. (That said, the argument in the preface suggests that the more incentivized player will win $\mathrm{BB}(\rho)$ at almost no running cost. But the absence of equilibria in the game gives neither this player nor her opponent any guidance as to how to play it.)
@@
There exist $C>0$ and $\eta,\kappa_{0}\in(0,1)$ such that, for $\kappa\in(0,\kappa_{0})$ and $\rho\in(0,1]$, $\big|\lambda_{\rm max}(\kappa,\rho)-1\big|\leq C\kappa^{\eta}$.
Remark. The result may be extended to the regime $\rho>1$ when $(\kappa,\rho)\in W$ (that is, $\kappa^{2}\rho\leq 1$), with $\eta=\eta(\rho)$ decaying to zero in the high-$\rho$ limit.
In the Epilogue, we report on $\lambda_{\rm max}:(0,1]^{2}\to[1,\infty)$ numerically, finding this function to have some remarkable features. The numerics prompt the following conjecture.
@@ 1.5.4. Structure of the article
The article is divided into two Parts.
Part I is devoted to the discrete Trail of Lost Pennies games. It develops an analytic framework for studying the ABMN($\kappa,\rho$) system, and classifies the games’ time-invariant Nash equilibria.
Part II returns to Brownian Boost. It presents a heuristic derivation and analytic study of the $\mathrm{BB}(\rho)$ ODE pair governing equilibrium behaviour, and then realises Brownian Boost as a fine-mesh, high-noise limit of the discrete games.
The Epilogue reports on numerical investigations, outlines several directions for further inquiry and is followed by a glossary of notation.
@@ 1.5.5. Acknowledgements
The author warmly thanks a referee of [24] whose pivotal role has been indicated in Section 1.2. He thanks several contributors in peer review whose helpful input has informed the structure and literature review of the article. He is supported by NSF-DMS grants $2153359$ and $2450726$.
Part I Discrete Tug-of-War Games
Chapter 2 Some basic symmetries and tools
The discrete Trail of Lost Pennies games are the object of study in this Part. Their analysis is organised around the $\mathbb{Z}$-indexed ABMN($\kappa,\rho$) system from Definition 1.2, which encodes equilibrium stakes and values. Central to our approach are structural features of this system, including the map $s:(0,\infty)\to(0,\infty)$ sending $\phi_{0}$ to $\phi_{1}$ introduced in Definition 1.8.
The aim of this three-chapter Part is to develop a rigorous analytic framework for the discrete games and to relate their equilibria to solutions of ABMN($\kappa,\rho$). In the present chapter, we assemble the basic tools needed for this programme. In Chapter 3, we show that $s$ acts as a shift on the $\phi$-sequence associated to any ABMN($\kappa,\rho$) element by Definition 1.20; we prove the explicit-form Theorem 1.11 for ABMN($\kappa,\rho$); and, by developing $s$-orbit asymptotics, we establish the fixed-parameter ABMN($\kappa,\rho$) Theorem 1.22. In Chapter 4, we complete the picture by proving the Nash–ABMN equivalence Theorem 1.3.
In this chapter, we focus on the foundational symmetries and constructions. In consecutive sections, we prove basic properties of $\phi_{0}$, $\phi_{1}$ and the resulting map $s(\phi_{0})=\phi_{1}$ and describe the escape behaviour of forward and backward $s$-orbits; introduce a role-reversal symmetry that inverts $s$; define the battlefield index of an element of ABMN($\kappa,\rho$) via its $s$-orbit; and introduce the Mina margin map, a key device for analysing equilibrium structure. In a final section, we analyse Penny Forfeit, the one-step sub-game of $\mathrm{TLP}(\kappa,\rho)$.
@@ 2.1. The bijections $\phi_{0}$ and $\phi_{1}$, and the orbit of $s$
The functions $\phi_{0}$ and $\phi_{1}$ are algebraically specified, in Definition 1.7. Our first result records some of their basic properties, which have permitted the specification of the map $s:(0,\infty)\to(0,\infty)$ that sends $\phi_{0}$ to $\phi_{1}$ in Definition 1.8.
@@
Figure 2.1. The main contours on the $(\kappa,\rho)$-map. The curve $C_{1}$ is the locus of $\kappa^{2}\rho=1$, which is the upper boundary of the region $W$ in (1.7); $C_{2}$ is the locus of $\kappa\rho=1+\sqrt{1-\kappa^{2}}$. The curves emanate from $(1,1)$, the point indexing the game studied in [24] and the upper-right corner of the unit box in which Nash-ABMN equivalence is established by Theorem 1.3. The $\mathrm{BB}(\rho)$-line $\kappa=0$ lies below $C_{1}$, which is indicative of how the ODE-pair Theorem 1.15 is valid for all $\rho\in(0,\infty)$.
@@
Writing $q$ for the right-hand denominator in (1.9), $\frac{\partial\,\phi_{0}(\kappa,\rho,\beta)}{\partial\beta}\,=\,P(\beta^{\rho})q^{-2}$ where of the coefficients of the quartic $P(x)=\sum_{i=0}^{4}h_{i}x^{i}$, $h_{0}=(1+\kappa)^{2}$ and $h_{4}=(1-\kappa)^{2}$ are evidently non-negative. That $h_{2}=2\big(3-\kappa^{2}(1+2\rho^{2})\big)$ is likewise follows from the conditions, which are weaker than our hypothesis, that $|\kappa|\leq 1$ and $\rho|\kappa|\leq 1$. The coefficients $h_{1}=4(1-\kappa)(1-\kappa\rho^{2})$ and $h_{3}=4(\kappa+1)(\rho^{2}\kappa+1)$ are also non-negative: in one case trivially; in the other, as our hypothesis is tailored to show; and with the sign of $\kappa$ determining which case applies.
@@
where $g(\kappa,x)=(1-\kappa)x^{2}+2(1-\rho\kappa)x+1+\kappa$. When $\kappa\in(0,1)$ and $\rho>0$, the quadratic $g(-\kappa,\bullet):\mathbb{R}\to\mathbb{R}$ is always positive, while for such $\kappa$-values, $g(\kappa,\bullet):\mathbb{R}\to\mathbb{R}$ is as well, because the discriminant sign condition $\rho<\kappa^{-1}(1+\sqrt{1-\kappa^{2}})$ has been hypothesised. Thus, $\phi_{0}$ exceeds $\phi_{1}$ when $\kappa\in(0,1)$.
@@
2.2. Role-reversal symmetry and the inverse of $s$
Recall the region $W$ of $(\kappa,\rho)$-parameter space from (1.7).
Proposition 2.5.
Let $(\kappa,\rho)\in W$. The function $s:(0,\infty)\to(0,\infty)$ from Definition 1.8(1) is invertible, with s1(x)= $s^{-1}(x)\,=\,1$/s(1/x) for  $x\in(0,\infty)$.
Proof. By definition, $s$ sends $\phi_{0}$ to $\phi_{1}$. Since both maps $\phi_{i}(\beta)=\phi_{i}(\kappa,\rho,\beta)$ are bijections $(0,\infty)\to(0,\infty)$ by Lemma 2.1(2), the inverse map $s^{-1}$ sending $\phi_{1}$ to $\phi_{0}$ is well defined. The formula s1(x)= $s^{-1}(x)\,=\,1$/s(1/x) amounts to $1/\phi_{0}=s(1/\phi_{1})$. To see this, note from (1.9) and (1.10) that when $\phi_{0}=\phi_{0}(\kappa,\rho,\beta)$ and $\phi_{1}=\phi_{1}(\kappa,\rho,\beta)$, we have that $1/\phi_{0}=\phi_{1}(\kappa,\rho,1/\beta)$ and $1/\phi_{1}=\phi_{0}(\kappa,\rho,1/\beta)$. So the sought equality $1/\phi_{0}=s(1/\phi_{1})$ is then the instance of $s(\phi_{0})=\phi_{1}$ corresponding to $1/\beta$. ∎
The preceding proof establishes the claimed identity by a direct algebraic calculation. It is conceptually helpful, however, to complement this argument with a game-theoretic point of view, which also motivates the term role reversal for the symmetry just identified. Some apparatus developed later in Part I is needed for this interpretation.
Two ingredients are required. First, the Nash–ABMN equivalence Theorem 1.3 furnishes a game-theoretic interpretation of any given $(a,b,m,n)\in$ ABMN($\kappa,\rho$). Second, while the proof above works with the formulaic definitions of $\phi_{0}$ and $\phi_{1}$, we will wish to invoke their interpretation in terms of sequence elements. In Lemma 3.3, we will establish the coincidence of these two descriptions, as discussed shortly after Definition 1.20: for $(a,b,m,n)\in$ ABMN($\kappa,\rho$), writing $\beta=\beta_{0}:=n_{1,-1}/m_{-1,1}$ yields $\phi_{0}=n_{0,-1}/m_{-1,0}$ and $\phi_{1}=n_{1,0}/m_{0,1}$.
With these identifications in place, the symmetry underlying Proposition 2.5 admits the following interpretation. Reflect gameplay governed by the strategy pair $(a,b)$ through the origin. The players now occupy opposite ends of the gameboard, and under the reflected dynamics each would be acting against her own interest. But their play makes sense if they now change ends. The resulting gameplay is governed by the strategy pair $(b(-\bullet),a(-\bullet))$, which extends naturally to
$\big(b(-\bullet),a(-\bullet),n(-\bullet),m(-\bullet)\big)\in\text{{ABMN}($\kappa,\rho$)}\,.$
Under this transformation, the parameters satisfy $\phi_{0}\mapsto 1/\phi_{1}$, $\phi_{1}\mapsto 1/\phi_{0}$, and $\beta_{0}\mapsto 1/\beta_{0}$, which explains both the formula $s^{-1}(x)=1/s(1/x)$ and the reciprocal $\beta$-parametrization that appears in the proof above.
The solution class ABMN($\kappa,\rho$) is invariant under $\mathbb{Z}$-shifts of the index, and the role-reversal symmetry identified here provides a second invariance of the system.
Corollary 2.6.
If $(a,b,m,n)$ is an element of ABMN($\kappa,\rho$), then so is
@@ 2.3. The battlefield index
Proof: (1). The function $s=s(\kappa,\rho,\bullet):(0,\infty)\to(0,\infty)$ belongs to $\rm IB$}, so it meets the decreasing map $1/x$ at exactly one $x\in(0,\infty)$.
@@ 2.4. The Mina margin map
In Definition 1.5, $\lambda_{\rm max}(\kappa,\rho)$ has been defined to be the supremum of the Mina margin $n_{\infty,-\infty}/m_{-\infty,\infty}$ over all ABMN($\kappa,\rho$) solutions. It is worth noting that the several symmetries enjoyed by ABMN($\kappa,\rho$) permit a more restricted supremum to be taken, and the Mina margin map is a useful device for making this point. Recall from Section 1.3.3 that for $x\in(0,\infty)$ there is a unique element of ABMN($\kappa,\rho$) with $m_{-\infty}=n_{\infty}=0$, $m_{\infty}=1$ and $\frac{n_{0,-1}}{m_{-1,0}}=x$. This is the standard solution $\big(a^{\rm st}_{i}(x),b^{\rm st}_{i}(x),m^{\rm st}_{i}(x),n^{\rm st}_{i}(x):i\in\mathbb{Z}\big)$.
@@
$\mathcal{M}_{\kappa,\rho}^{-k,k}(x)\,\longrightarrow\,\mathcal{M}_{\kappa,\rho}(x)\,$, (2.1)
@@
As $k\to\infty$, the left-hand side converges to $\mathcal{M}_{\kappa,\rho}\big(s(x)\big)$, by (2.1) with $x\to s(x)$, while the right-hand side converges to $\frac{n^{\rm st}_{\infty,-\infty}}{m^{\rm st}_{-\infty,\infty}}=\mathcal{M}_{\kappa,\rho}(x)$ by (2.1) and the decay of high-indexed $m$- and $n$-differences in Theorem 1.22. Hence $\mathcal{M}_{\kappa,\rho}(s(x))=\mathcal{M}_{\kappa,\rho}(x)$ for $x\in(0,\infty)$.
@@
(2). There is a Mina margin map associated to each finite trail. For the gameboard $\llbracket-k,k\rrbracket$, it takes the form